A fully charged parallel-plate capacitor remains connected to a battery while you slide a dielectric between the plates. Do the following quantities increase, decrease or remain the same? (i) C (ii) Q (iii) Change in V (iv) the energy stored in the capacitor
1
Expert's answer
2020-11-05T07:32:34-0500

We have "C = \epsilon_0\times A/d" for free space.


Where "\epsilon_0" is the permittivity of free space.

"A" is the area of parallel plate capacitor

"d" is the distance between parallel plates


When a dielectric with dielectric constant K is inserted the formulae becomes


"C = K \times\epsilon_0\times A/d"


(i) Hence from above equation The capacitance increases by K times.


(ii) We have Charge "Q = C \times V" Where "Q" is the charge and "V" is the voltage across capacitor.


As capacitance increases by K times Charge increases by K times.


(iii) As battery is kept connected There is no change in V i.e change in V =0.


(iv) = Energy = "1/2 \times C \times V^2" This too is directly proportional to Capacitance. That means

Energy increases by K times